Carbon-14 is a radioactive element with a half-life of about 5,730 years. Carbon-14 is said to decay exponentially. The decay rate is 0.000121. We start with one gram of carbon-14. We are interested in the time (years) it takes to decay carbon-14. (A) Find the value k such that P(x < k) = 0.6. (Round your answer to two decimal places.)
Answer #1

The require probability is defined by cumulative distribution function which is,

P(x<k)=1-e^{-\lambda k}

Where,

P(x < k) = 0.6

λ 0.000121

Hence,

P(x<k)=1-e^{-\lambda k}=0.6\Rightarrow e^{-\lambda k}=0.4

In(2.5) 0.4 taking natural logk

In(2.5 0.916291 -.- 0.000121

k 7572.65068 years